What is the real increase in value if $2,500 is invested for one year at 7% interest and the rate of inflation during that time is 2.1%
Answer #1
Nominal rate of interest 7%
Inflation rate 2.10%
Real rate of return = (1+Nominal rate) /(1+Inflation rate) -1
1.07 /1.021 - 1 =4.80%
Amount invested = 2500
Amount after one year = Amount invested (1+reat rate)
2500 *1.048 = 2620
Real increase in value = 2620-2500 =120